ABBA Corp has issued ABBA Bond on discount on April 6, 2010. This 8 percent bond is

redeemable at par on April 6, 2030, or on any optional redemption date. The expected

yield of the bond on the date of issue was 10 percent p.a. Coupon for the bond is paid

quarterly. The earliest possible ABBA Corp could redeem the whole bond (but not in part)

is at the end of its 12-year deferment period with a 12 percent call premium. An investor

has decided to purchase this bond today, December 21, 2019, when the interest rate has

increased by 2 percent (from the last known rate).

i) Compute his current yield on the date of purchase.

ii) Compute his yield to call.


1
Expert's answer
2022-01-10T17:34:03-0500

i)The nominal value will be 1000, and the purchase price will be 900

"i=\frac{N-P}{P}\times\frac{365}{n}\times100"

N=1000

P=900

n=3546( 06.04.2010-21.12.2019)

"i=\frac{1000-900}{900}\times\frac{365}{3546}\times100=1.14"


ii)"YTC=\frac{C+\frac{Pc-P}{n}}{\frac{Pc+P}{2}}"

"C=\frac{0.08}{4}\times1000=20"

"n=4\times12=48"


Let the trade price be 950

Pc=950 P=900


"YTC=\frac{C+\frac{Pc-P}{n}}{\frac{Pc+P}{2}}=\frac{20+\frac{950-900}{48}}{\frac{950+900}{2}}=0.0227" or 2.27%